Find the area of a triangle whose sides have length 6,7, and 10.
Answer by checkley71(8403)

SEEING AS THIS IS NOT A RIGHT TRIANGLE WE USE THE FORMULA:
A=SQRT[S(S-A)(S-B)(S-C)]WHERE S=THE SUM OF ALL SIDES DIVIDED BY 2 OR:
(6+7+10)/2=23/2=11.5
A=SQRT[11.5(11.5-6)(11.5-7)(11.5-10)]
A=SQRT[11.5*5.5*4.5*1.5]
A=SQRT426.9375
A=20.66 ANSWER.